OpenStudy (anonymous):

For the Function f(x)=2X^3-7x^2+4x+1 a) find the instantaneous rate of change at x=0 and x=1 b)Is the function increasing or decreasing at x=0? x=1? c)Do you expect a local min or max when 0

OpenStudy (anonymous):

f'(x) = 6x^2 - 14x + 4 f'(1) = 6(1)^2 - 14(1) + 4 f'(1) = 6 - 14 + 4 f'(1) = -4 f'(0) = 6(0)^2 - 14(0) + 4 f'(0) = 4 since the instantaneous rate of change (read: derivative) at x = 0 is positive, the function is increasing at x = 0 since the instantaneous rate of change (read: derivative) at x = 1 is negative, the function is decreasing at x = 1 since from left to right (the graph of x = 0 to x = 1), the function switches from increasing to decreasing, it looks like there would be a maximum between 0 and 1

OpenStudy (anonymous):

Refer to the attached Mathematica solution.